Jaunice is redecorating her room. She wants to paint a design on a
rectangular section of her wall that is at least 76.5 square feet. The height of the design will be 8.5 feet. What is the minimum length of the design Juanice will paint on the wall?
A. 8 feet
B. 8.5 feet
C. 9 feet
D. 9.5 feet

Answer by Boreal(15235)

area is l*h=8.5*l=76.5
l=9 feet
C